OpenStudy (anonymous):

simplify (2 + 3.2 x 15) x (8 - 2 to the second power

OpenStudy (anonymous):

some1 please help

OpenStudy (anonymous):

okay (2+3.2x15)=50 (8-2)^2=36 36x50=1800

OpenStudy (anonymous):

Does that help? :)

OpenStudy (anonymous):

yes thank you so much

OpenStudy (anonymous):

no problem and thank you for the metal :)!!